AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
DESPEC
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Nov 7 14:46:36 2025
Author
*
:
Subject
*
:
- Ch 2 and 3 of CAEN HV module changed to negative polarity, Ch 0,1 still with positive polarity - Bias changed from the n+n to the p+n side (now Ch 2,3) - att. 1 - EXPERIMENTS/AIDA/2025Oct28-14.55.15 -> Slow comparitor -> 0x64 (1MeV) p+n Pos. bias & n+n Neg. bias - Attachment 7: aida01 had its threshold lowered to 500 keV -> Rate increased from 500 to 6500 Hz - Pulser inputs plugged in. We have input a 1V pulser input into the data and have measured a peak in aida01 with a FWHM = 75 channel see attachment 9 and a FWHM in aida 14 (DSSD2) of also 76 channels which is similar to the results obtained in June from Nabil. We cannot see anything from the n+n side as the signals are too noisy. Probably best to go with a higher trigger rate on the pulser and check with log scale tomorrow. It might also be worth checking the grounding on the y-strips as the problem with biasing with positive polarity is still seen in aida02 and aida04 (high noise repetition0 signal, attachment 2). Conclusion is that we have reproduced in part the result from Nabil. Tomorrow we will continue with the tests by lowering the thresholds and hopefully collect some data :-)
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898